OpenStudy (anonymous):

Which value of x makes the following equation true? 2(x + 10) = x + 24 A. 24 B. 14 C. 4 D. -6

OpenStudy (anonymous):

Please help me

OpenStudy (anonymous):

ur answer is C.4

OpenStudy (anonymous):

can i please have a medeal and get help on a ?

OpenStudy (anonymous):

I gave you a medal and I can try and help you

OpenStudy (anonymous):

thnxs and okay

OpenStudy (anonymous):

atag me in it

OpenStudy (anonymous):

tag*

OpenStudy (anonymous):

*PurplePanda312 -(-21+8)=21+(-8) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

'@PurplePanda312 '

OpenStudy (anonymous):

'@PurplePanda312

OpenStudy (anonymous):

I don't know. Sorry I can't help you.

OpenStudy (anonymous):

okay can u try this one Which is an example of the opposite of a sum property? A. –17 + 32 = 32 + (–17) B. –26 + (4 + 8) = (–26 + 4) + 8 C. –5(6 + 3) = –5(6) + –5(3) D. –(–17 + 5) = 17 + (–5)